# SketchMind: A Multi-Agent Cognitive Framework for Assessing Student-Drawn Scientific Sketches

### v1

@@ 3.1 Cognitive Framework: Bloom’s Taxonomy in Sketch Understanding
Bloom’s Taxonomy defines six ascending levels of cognitive complexity:
=$\mathcal{B}=\{\mathrm{Remember},\mathrm{Understand},\mathrm{Apply},\mathrm{Analyze},\mathrm{Evaluate},\mathrm{Create}\}.$ (1)
In SketchMind, scientific sketches are not simply visual representations but are treated as cognitive artifacts that externalize internal mental models. Each node and edge in an SRG is assigned a Bloom level to represent the depth of conceptual engagement exhibited by the student. It is worth noting that different assessments may have different highest Bloom’s level. We associate ascending ordered number to the Bloom’s level such as 1 for Remeber and 6 for Create. These assignments are used both for semantic evaluation and for guiding personalized feedback.
@@ 4 Experimental Setup
We evaluate SketchMind using two distinct model configurations to comprehensively examine its performance and versatility across closed-source and open-source MLLM frameworks. In Configuration 1, we utilize GPT-4o, GPT-4.1, GPT-4.1-nano, o3, and o4-mini (all latest GPT models with different reasoning capabilities for comprehensive evaluation) accessed via the OpenAI API. This setup leverages black-box prompting tailored explicitly to encourage structured JSON output and support multimodal reasoning. Configuration 2 deploys open-source MLLMs, specifically the FP8-quantized Llama-4 Maverick and Scout (400B and 109B parameters, with 17B active parameters), running locally on two NVIDIA H100 GPUs, facilitated by Hugging Face Transformers (version 4.39+) and PyTorch (version 2.1+). We keep the same model for each agent to maintain the model performance consistency, for example, if targeted GPT-4o, then each agent uses GPT-4o and mainatin local chat session.
For closed-source configurations, inference requests are systematically routed through the OpenAI API, meticulously logging timestamps, model versions, and full prompt-response interactions to ensure transparency and reproducibility. Open-source models are executed on-premises, utilizing dual H100 GPUs with CUDA 12.1 and cuDNN 8.9, loading quantized model weights from the Hugging Face Model Hub, and multimodal data inputs are managed by the Llama4Processor.
@@
Given this gap, we base our study on a rigorously developed dataset originally introduced by Zhai et al. [28], which has since become one of the most appropriate and widely recognized resources for evaluating automated reasoning over student-generated scientific models. This dataset was adapted from the NGSA (Next Generation Science Assessment) initiative [10], which aligns tightly with the NGSS framework [24] and has been used by AIED researchers to assess students’ conceptual understanding through multimodal evidence [15, 14]. Comprehensive details about dataset selection rationale and statistics can be seen in Appendix. LABEL:sec:appendix:datase_stats.

### v2

@@ 3.1 Cognitive Framework: Bloom’s Taxonomy in Sketch Understanding
Bloom’s Taxonomy structures learning objectives into six ascending levels of cognitive complexity [6]:
$\mathcal{B}=\{\mathrm{Remember},\mathrm{Understand},\mathrm{Apply},\mathrm{Analyze},\mathrm{Evaluate},\mathrm{Create}\}.$ (1)
These range from basic recall of knowledge (Remember) to the synthesis of novel ideas (Create). This cognitive hierarchy has long served as a foundation in science education for designing assessments and scaffolding learning [13]. Prior work further shows that aligning instructional technologies with Bloom’s levels supports measurable gains in higher-order thinking [8]. In SketchMind, scientific sketches are not merely visual representations but are conceptualized as cognitive artifacts that externalize learners’ mental models. To operationalize this, we annotate each node (concept) and edge (relation) in the Sketch Response Graph (SRG) with a Bloom level, a process we call Bloom-Level Annotation. This provides a fine-grained measure of the depth of conceptual engagement demonstrated by the student.
To implement this systematically, rubric statements are parsed to extract key verbs and criteria, which are then mapped to Bloom levels using a curated lexicon [24]. Ambiguous cases are resolved with a fine-tuned classifier. The resulting numeric levels (1 for Remember through 6 for Create) are encoded as attributes on SRG nodes and edges. These attributes directly support semantic similarity scoring, nuanced evaluation, and feedback generation.
Pedagogical Integration.
Each gold-standard SRG is labeled at the level of its highest Bloom demand. Student sketches are then evaluated by comparing the Bloom-level annotations of their SRGs against this reference. This supports both diagnostics (identifying the highest level achieved) and adaptive feedback. For example, if a student’s sketch demonstrates Understand, the system can generate targeted prompts or visual hints nudging them toward Apply. Such progression-oriented scaffolding is consistent with educational research showing that adaptive support aligned with Bloom’s hierarchy fosters deeper learning in STEM domains [13, 8].
A schematic illustration (see Fig. 1) depicts this process: a rubric statement (e.g., “Describe the direction of dye movement”) maps to an SRG node (Dye Direction), which is then assigned a Bloom level (Understand). This visual link between task, rubric, SRG structure, and Bloom’s hierarchy enhances transparency for both learners and instructors.
@@ 4 Experimental Setup
We evaluate SketchMind using two distinct model configurations to comprehensively examine its performance and versatility across closed-source and open-source MLLM frameworks. In Configuration 1, we utilize GPT-4o, GPT-4.1, GPT-4.1-nano, o3, and o4-mini (all latest GPT models with different reasoning capabilities for comprehensive evaluation) accessed via the OpenAI API. This setup leverages black-box prompting tailored explicitly to encourage structured JSON output and support multimodal reasoning. Configuration 2 deploys open-source MLLMs, specifically the INT4-quantized Llama-4 Maverick and INT8-quantized Scout (400B and 109B parameters respectively, with 17B active parameters), running locally on four NVIDIA H100 GPUs, facilitated by Hugging Face Transformers (version 4.39+) and PyTorch (version 2.1+). We keep the same model for each agent to maintain the model performance consistency, for example, if targeted GPT-4o, then each agent uses GPT-4o and mainatin local chat session.
For closed-source configurations, inference requests are systematically routed through the OpenAI API, meticulously logging timestamps, model versions, and full prompt-response interactions to ensure transparency and reproducibility. Open-source models are executed on-premises, utilizing four H100 GPUs with CUDA 12.1 and cuDNN 8.9, loading quantized model weights from the Hugging Face Model Hub, and multimodal data inputs are managed by the Llama4Processor.
@@
Given this gap, we base our study on a rigorously developed dataset originally introduced by Zhai et al. [30], which has since become one of the most widely recognized resources for evaluating automated reasoning over student-generated scientific models. This dataset, adapted from the NGSA (Next Generation Science Assessment) initiative [10], aligns closely with the NGSS framework [26] and has been used by AIED researchers to assess students’ conceptual understanding through multimodal evidence [16, 15]. Comprehensive details about dataset selection rationale and statistics are provided in Appendix LABEL:sec:appendix:datase_stats. To ensure pedagogical validity, we conducted a structured human evaluation of model-generated feedback. Four domain-expert educators, each with graduate-level training in science education, independently assessed the pedagogical quality of system responses. Each rater evaluated a stratified random sample of 890 student-generated sketches (25% of the dataset), ensuring balanced representation across baseline models (GPT-4o, GPT-4.1, O3), grade levels, and science task types.
All evaluators participated in a calibration phase: they jointly annotated and discussed 10 representative examples to align rubric interpretation, and then independently scored the same 10 sketches. Inter-rater reliability, measured using Quadratic Weighted Kappa, reached $\kappa=0.83$, which is consistent with benchmarks for expert judgment in science assessment [11]. For the main study, each response was rated by at least two experts in a double-blind manner.
Evaluation guidelines were based on an extended rubric adapted from Zhai et al. [30] and tailored for formative science assessment. Experts rated each response along three dimensions: 1) Clarity: Whether the textual and visual feedback was understandable and actionable. 2) Conceptual Accuracy: Whether suggested modifications and explanations correctly reflected the target scientific concepts. And 3)Instructional Value: The feedback’s potential to promote student learning and progression along Bloom’s taxonomy. Each dimension was scored on a 5-point ordinal scale (1–5). Consistency across raters was supported through the initial calibration process, and ratings were conducted under double-blind conditions to mitigate bias. Framework–human agreement rates, along with detailed rubric codebooks and annotated examples, are reported in Section 4 and the supplementary material.